# Quantifying sleep apnea heterogeneity using hierarchical Bayesian modeling

### v1

@@ 3.3.1 Point estimation
where each cluster center $\overline{\theta}_{k}$ is a random variable that varies over posterior samples. Note that (1) and (2) are not equivalent and, in general, the learned partitions defined by $\hat{c}$ and $\hat{c}^{\prime}$ will differ. In the results that follow, we focus on the formulation (1); we are interested in interpreting the cluster summaries in the context of the estimated cluster centers, so it is appealing to obtain a principled point estimate of these centers jointly with the assignments. If we first compute $\hat{c}^{\prime}$ using (2) and then summarize the centers $\hat{b}^{\prime}$ as posterior means conditional on $\hat{c}^{\prime}$, we produce a suboptimal solution $(\hat{c}^{\prime},\hat{b}^{\prime})$ with respect to (1). However, we also implement formulation (2) for comparison and summarize the results in Supplementary Information.
@@ 5.1 Exploratory data analysis
Before fitting our model, we performed some data visualization to get a sense of the heterogeneity in OSA event occurrence in the APPLES participants. Figure 1 shows the distribution of AHI values, both in general, as used when classifying OSA severity, and specifically during REM and non-REM sleep. Observe that the average AHI appears to be somewhat higher during REM sleep compared to non-REM sleep.
Figure 1: Distribution of REM, non-REM, and overall AHI values for patients in the APPLES study.
To provide additional context to the AHI distributions, Figure 2 shows the distributions of the average event durations at the patient level during non-REM and REM sleep. Here again, the event durations appear longer on average in REM sleep compared to non-REM sleep. This not only adds to the evidence that OSA intensity may be higher for the average person in REM sleep compared to non-REM, but actually suggests that the difference between AHI values in REM vs. non-REM may in some sense understate the higher rate during REM sleep, since longer average events lead to a lower AHI given the same distribution of times between events. (E.g., a patient with two 10-second events each minute will have double the AHI of a patient with one 40-second event each minute, even though they both have an average inter-event time of 20 seconds, and more of the second patient’s sleep time is spent having events.) This suggests our approach of modeling inter-event times as the time between the end of one event and the start of the next may more reasonably capture the difference in in OSA intensity compared to AHI.
Figure 2: Distribution of average hypopnea/apnea event durations for patients in the APPLES study. Four patients are omitted for REM sleep: two with zero events during REM, and two with outlier mean event lengths of 219 and 188 seconds.
Finally, Figure 3 shows the distributions of time spent in REM vs. non-REM sleep in the sample of patients. Clearly, the average patient spends far more time in non-REM than in REM. This suggests that sleep stage-specific event rates can affect sleep quality and health through different pathways. On the one hand, since non-REM represents the majority of sleep, an elevated rate of non-REM events has a larger impact on the total number of events in a night than an elevated REM rate, as reflected by the similarity in the non-REM and overall AHI distributions in Figure 1. Thus, one can imagine that the rate during non-REM could be particularly impactful on certain health outcomes of interest. However, since REM sleep is believed to be particularly important for learning and memory (Peever and Fuller, 2017), if an elevated rate of REM events causes a patient’s small but important amount of REM sleep to be interrupted or cut short, this could also be highly detrimental.
Figure 3: Distributions of time in non-REM sleep, time in REM sleep, and total time asleep for patients in the APPLES study.
@@ 5.3.1 Model parameters
Figures 5-6 summarize the posterior distributions for the transition and inter-event model parameters. In the left panel of Figure 5, observe that the posteriors of all $\mu$ parameters are strongly negative, indicating that for both REM and non-REM sleep, patients are overwhelmingly likely to stay in the same stage from one epoch to the next. The 95% posterior credible intervals for $\tau_{RN}$, $\tau_{RA}$, $\tau_{NR}$, and $\tau_{NA}$ are all strictly positive, indicating that having an apnea event in a given epoch makes the average patient more likely to transition out of their current sleep stage. Note that $\tau_{RA}$ and $\tau_{NA}$ have magnitudes larger than $\tau_{RN}$ and $\tau_{NR}$, respectively, indicating that apnea events have a particularly strong effect on the probability of awakening. In the right panel of Figure 5, note that the baseline transition parameters $\gamma$ appear to have a larger variance for transitions out of REM than for transitions out of non-REM, while the opposite is true for the parameters $\alpha$. This suggests that there may be more baseline patient variation in transitions from REM sleep, but more variation in the effect of OSA events during non-REM sleep.
Figure 5: Posterior summaries of fixed effects (left) and random effect variances (right) from the stage transition Markov model. Dots represent posterior means, and intervals represent 95% posterior credible intervals.
As expected from our EDA in Section 5.1, the left panel of Figure 6 shows that the baseline rate of events is substantially higher during REM sleep compared to non-REM. The random effect variances shown in the right panel of Figure 6 are very similar between REM and non-REM, with the posterior mean slighlty higher for non-REM. However, since these random effects $\phi_{ik}$ affect patient-level expected interevent times multiplicatively, the actual patient-level variability is higher during REM sleep than non-REM, due to the higher baseline rate during REM. Finally, note that given the model setup from Section 3.1, the variances observed for the $\phi_{ik}$s indicate substantial patient variation in interevent times. If we take each of the $\phi_{ik}$ variances to be about $0.7$, this corresponds to a patient with a rate one standard deviation above the mean having an expected interevent time over five times shorter than a patient whose rate is one standard deviation below the mean.
Figure 6: Posterior summaries of fixed effects (left) and random effect variances (right) from the apnea and hypopnea inter-event time model. Dots represent posterior means, and intervals represent 95% posterior credible intervals.
@@
Figure 8 shows the posterior means of the factor loadings in the matrix $\Lambda$, aligned for interpretability using the method of Poworoznek et al. (2021). In general, the interpretations of these factors yield similar insights to examining the correlations in Figure 7. Factors 1 and 5 yield the positive correlations among the REM transition parameters, factor 4 yields the positive correlation between $\phi_{R}$ and $\phi_{N}$, and factors 2 and 3 are primarily responsible for the negative correlations discussed above.
Figure 8: Posterior mean factor loadings after alignment with the procedure of Poworoznek et al. (2021). Loadings for which the $95\%$ credible interval contains zero have been set to zero.
@@ 5.3.2 Clustering
The cluster assignment point estimates are shown in Figure 9, indicated by colored points showing each $\theta_{i}$ projected into two dimensions. The vertical axis is simply $\phi_{R}$, while the horizontal axis is the first principal component score for PCA performed on the posterior means of $(\gamma_{RA},\gamma_{NA},\alpha_{RA},\alpha_{NA},\gamma_{RN},\gamma_{NR},\alpha_{RN},\alpha_{NR})$. The vector of the first associated principal component is $(0.28,0.06,0.48,0.23,0.33,0.50,0.52,-0.07)$, suggesting that a positive value of pc1 roughly corresponds to more disruption of the sleep stages, both at baseline and due to OSA events. Given this interpretation, the projected locations of the clusters appear to correspond to the above interpretation of their centers.
Figure 9: Bayes-optimal k-means point estimate of cluster assignments based on random effect vectors $\theta_{i}$. pc1 refers to the first principal component vector computed with the posterior means of the dynamics model random effects.
To illustrate the uncertainty quantification produced by equation (3), Figure 10 shows the probabilities of assignment to cluster 1. Although the majority of patients have probabilities very close to 0 or 1, a number of observations near the boundary with other clusters have meaningful levels of uncertainty in their assignment.
Figure 10: Posterior probability of assignment to cluster 1, computed with equation (3).
@@ 5.4 Relationship of cluster assignment with other variables
To illustrate that random effects in our model and the clusters they produce can generate insights missed by a simple AHI approach, Table 5 shows the results of two linear regressions. The first regresses the BSRT score on the cluster label, age, and sex, while the second regresses BSRT score on AHI, age, and sex. Sex and age are found to be significant in both models, with higher age and being male associated with poorer performance. In model 1, cluster 1 has a significantly better score than reference cluster 4. However, despite the variation in AHI between the groups and the fact that group 4 has by far the highest AHI, in model 2 the coefficient for AHI is not significant at a significance level of 0.05, suggesting that there may be more to the story when relating OSA severity to a cognitive outcome like BSRT. One aspect of this is that while cluster 1 does not have the lowest AHI, it does have the lowest average $\phi_{R}$, suggesting that the rate of events may matter more in REM sleep than non-REM. Note that group 3 has the lowest AHI and the lowest average $\phi_{N}$, but does not have significantly better performance than group 4 in model 1. Similarly, while cluster 2 also has substantially lower AHI than cluster 4, it also does not differ significantly in BSRT, and its coefficient has a negative point estimate. One interpretation of these results is that the higher AHI of cluster 4 is balanced by cluster 2’s more highly disrupted sleep, which suggests that examining OSA event occurrences and sleep dynamics together as our model does is important for understanding drivers of symptom severity.
@@ 5.5 Alternate clustering based on simple summary statistics
Finally, in Table 7, we examine the differences in the BSRT scores among these alternate groups. Setting cluster 4 as the reference, we find that after controlling for sex and age, patients in cluster 1 have a significantly higher expected BSRT score at a significance level of $0.05$. Taken together, the above results suggest that in practice, simple summary statistics such as time in each sleep stage and stage-specific AHI can help give complementary insights to those generated by our model. More broadly, examining how other variables differ by the clusters generated by our model’s random effects can help reveal other variables that could be key to better understanding sleep differences among OSA patients.
[…]
Regression 3 Intercept 62.11 1.36 45.81 0.00
cluster1 1.79 0.80 2.23 0.03
cluster2 1.74 0.91 1.92 0.06
cluster3 0.63 0.84 0.76 0.45
age -0.21 0.02 -9.75 0.00
sexMale -3.94 0.58 -6.84 0.00
Table 7: Results for a linear regression of BSRT score on alternate cluster labels with reference level 4, after accounting for age and sex.
@@ 6 Discussion
In this work, we investigated the heterogeneity in sleep stage dynamics and stage-specific rates of apnea and hypopnea events in OSA patients by analyzing data from the APPLES study, which generated several novel insights. We found that for the average patient, having an OSA event during a given epoch increases the probability of awakening or changing sleep stages in the following epoch, although both the baseline probabilities and the magnitude of this change vary substantially among patients. Also, for the average patient, the expected wait time between OSA events is substantially shorter during REM sleep than during non-REM, although again there is significant patient variation in rates. By modeling the covariance of our model’s random effects, we revealed that during REM sleep, patient baseline rates of transitions to Awake or non-REM are positively correlated with one another, as well as being correlated with greater impact of OSA events on these transitions, while during non-REM sleep, baseline rates of transitions to REM or Awake are negatively correlated with the corresponding shifts due to events. Finally, we clustered patients into four interpretable groups based on their random effects. This revealed that other demographic and health variables, such as age, sex, and BMI, vary significantly by event rates and sleep dynamics. Turning to the cognitive performance measure BSRT, we determined that relative to group 4, which had the highest average AHI, cluster 1 had significantly better BSRT scores after correcting for age and sex. This result was noteworthy because, while cluster 1 had the lowest rate of events during REM sleep, it did not have the lowest overall AHI, supporting the hypothesis that events in REM sleep may be particularly harmful to cognitive outcomes. In contrast to our clustering approach, a regression of BSRT on AHI directly did not yield a significant association after correcting for sex and age, further emphasizing the limitations of an AHI-based analysis.
@@
7 Appendix
7.1 Proof of Proposition 3.1
Observe that, letting all expectations be over the posterior distribution $\pi_{post}$, we have
$\left(\hat{c},\hat{b}\right)$ $=\operatorname*{arg\,min}_{c:|C|=K,\,\,b\in\mathbb{R}^{K\times d}}\sum_{k=1}^{K}\sum_{i\in C_{k}}\mathbb{E}\left[||\theta_{i}-b_{k}||^{2}\right]$
$=\operatorname*{arg\,min}_{c:|C|=K,\,\,b\in\mathbb{R}^{K\times d}}\sum_{k=1}^{K}\sum_{i\in C_{k}}\mathbb{E}\left[||(\theta_{i}-\mathbb{E}[\theta_{i}])+(\mathbb{E}[\theta_{i}]-b_{k})||^{2}\right]$
$=\operatorname*{arg\,min}_{c:|C|=K,\,\,b\in\mathbb{R}^{K\times d}}\sum_{k=1}^{K}\sum_{i\in C_{k}}\mathbb{E}[||\theta_{i}-\mathbb{E}[\theta_{i}]||^{2}]+\mathbb{E}[(\theta_{i}-\mathbb{E}[\theta_{i}])^{T}(\mathbb{E}[\theta_{i}]-b_{k})]+\mathbb{E}[||\mathbb{E}[\theta_{i}]-b_{k}||^{2}]$
$=\operatorname*{arg\,min}_{c:|C|=K,\,\,b\in\mathbb{R}^{K\times d}}\sum_{k=1}^{K}\sum_{i\in C_{k}}||\mathbb{E}[\theta_{i}]-b_{k}||^{2}$
where we have used that in the second-to-last line, the first term does not depend on $(c,b)$, the second term is equal to zero, and the outer expectation in the third term can be dropped. $\square$

### v2

@@ 3.3.1 Point estimation
where each cluster center $\overline{\theta}_{k}$ is a random variable that varies over posterior samples. Note that (1) and (S1) are not equivalent and, in general, the learned partitions defined by $\hat{c}$ and $\hat{c}^{\prime}$ will differ. In the results that follow, we focus on the formulation (1); we are interested in interpreting the cluster summaries in the context of the estimated cluster centers, so it is appealing to obtain a principled point estimate of these centers jointly with the assignments. If we first compute $\hat{c}^{\prime}$ using (S1) and then summarize the centers $\hat{b}^{\prime}$ as posterior means conditional on $\hat{c}^{\prime}$, we produce a suboptimal solution $(\hat{c}^{\prime},\hat{b}^{\prime})$ with respect to (1). However, we also implement formulation (S1) for comparison and summarize the results in Supplementary Information.
@@ 5.1 Exploratory data analysis
Before fitting our model, we performed some data exploration, summarized in Table 3. Observe that average AHI appears to be somewhat higher during REM sleep compared to non-REM sleep. Interestingly, event durations also appear to be longer in REM vs. non-REM. This not only adds to the evidence that OSA intensity may be higher for the average person in REM sleep compared to non-REM, but actually suggests that the difference between AHI values in REM vs. non-REM may in some sense understate the higher rate during REM sleep, since longer average events lead to a lower AHI given the same distribution of times between events. (E.g., a patient with two 10-second events each minute will have double the AHI of a patient with one 40-second event each minute, even though they both have an average inter-event time of 20 seconds, and more of the second patient’s sleep time is spent having events.) This suggests our approach of modeling inter-event times as the time between the end of one event and the start of the next may more reasonably capture the difference in in OSA intensity compared to AHI.
Finally, the average patient spends far more time in REM than non-REM sleep. This suggests that sleep stage-specific event rates can affect sleep quality and health through different pathways. On one hand, since non-REM represents the majority of sleep, an elevated rate of non-REM events has a larger impact on the total number of events in a night than an elevated REM rate. However, since REM sleep is believed to be particularly important for learning and memory (Peever and Fuller, 2017), if an elevated rate of REM events causes a patient’s small but important amount of REM sleep to be interrupted or cut short, this could also be highly detrimental. Box plots showing distributions of the quantities in Table 3 are included in the Supplementary Information.
AHI Mean event duration (seconds) Time spent in stage (hours)
REM 49.6 (22.7) 27.3 (11.4) 1.1 (0.5)
non-REM 43.5 (24.4) 24.1 (6.7) 5.2 (0.9)
Table 3: Patient means and standard deviations for stage-specific AHI, event durations, and time spent in each sleep stage.
@@ 5.3.1 Model parameters
Figure 2 summarizes the posterior distributions for the transition and inter-event model parameters. In the top left panel, observe that the posteriors of all $\mu$ parameters are strongly negative, indicating that for both REM and non-REM sleep, patients are overwhelmingly likely to stay in the same stage from one epoch to the next. The 95% posterior credible intervals for $\tau_{RN}$, $\tau_{RA}$, $\tau_{NR}$, and $\tau_{NA}$ are all strictly positive, indicating that having an apnea event in a given epoch makes the average patient more likely to transition out of their current sleep stage. Note that $\tau_{RA}$ and $\tau_{NA}$ have magnitudes larger than $\tau_{RN}$ and $\tau_{NR}$, respectively, indicating that apnea events have a particularly strong effect on the probability of awakening. In the top right panel of Figure 2, note that the baseline transition parameters $\gamma$ appear to have a larger variance for transitions out of REM than for transitions out of non-REM, while the opposite is true for the parameters $\alpha$. This suggests that there may be more baseline patient variation in transitions from REM sleep, but more variation in the effect of OSA events during non-REM sleep.
Figure 2: Posterior summaries of the Markov model fixed effects (top left) and random effect variances (top right), and the apnea and hypopnea inter-event time model fixed effects (bottom left) and random effect variances (bottom right). Dots represent posterior means, and intervals represent 95% posterior credible intervals.
As expected from our EDA in Section 5.1, the bottom left panel of Figure 2 shows that the baseline rate of events is substantially higher during REM sleep compared to non-REM. The random effect variances shown in the bottom right panel are very similar between REM and non-REM, with the posterior mean slightly higher for non-REM. However, since these random effects $\phi_{ik}$ affect patient-level expected interevent times multiplicatively, the actual patient-level variability is higher during REM sleep than non-REM, due to the higher baseline rate during REM. Finally, note that given the model setup from Section 3.1, the variances observed for the $\phi_{ik}$s indicate substantial patient variation in interevent times. If we take each of the $\phi_{ik}$ variances to be about $0.7$, this corresponds to a patient with a rate one standard deviation above the mean having an expected interevent time over five times shorter than a patient whose rate is one standard deviation below the mean.
@@
In the online supplement, we summarize the posterior means of the factor loadings in the matrix $\Lambda$, aligned for interpretability using the method of Poworoznek et al. (2021). In general, the interpretations of these factors yield similar insights to examining the correlations in Figure 3. Factors 1 and 5 yield the positive correlations among the REM transition parameters, factor 4 yields the positive correlation between $\phi_{R}$ and $\phi_{N}$, and factors 2 and 3 are primarily responsible for the negative correlations discussed above.
@@ 5.3.2 Clustering
The cluster assignment point estimates are shown in the left panel of Figure 4, indicated by colored points showing each $\theta_{i}$ projected into two dimensions. The vertical axis is simply $\phi_{R}$, while the horizontal axis is the first principal component score for PCA performed on the posterior means of $(\gamma_{RA},\gamma_{NA},\alpha_{RA},\alpha_{NA},\gamma_{RN},\gamma_{NR},\alpha_{RN},\alpha_{NR})$. The vector of the first associated principal component is $(0.28,0.06,0.48,0.23,0.33,0.50,0.52,-0.07)$, suggesting that a positive value of pc1 roughly corresponds to more disruption of the sleep stages, both at baseline and due to OSA events. Given this interpretation, the projected locations of the clusters appear to correspond to the above interpretation of their centers.
Figure 4: Bayes-optimal K-means point estimate of cluster assignments based on random effect vectors $\theta_{i}$ (left), and posterior probability of assignment to cluster 1, computed with equation (3) (right). pc1 refers to the first principal component vector computed with the posterior means of the dynamics model random effects.
To illustrate the uncertainty quantification produced by equation (3), the right panel of Figure 4 shows the probabilities of assignment to cluster 1. Although the majority of patients have probabilities very close to 0 or 1, a number of observations near the boundary with other clusters have meaningful levels of uncertainty in their assignment.
@@ 5.4 Relationship of cluster assignment with other variables
To illustrate that random effects in our model and the clusters they produce can generate insights missed by a simple AHI approach, Table 6 shows the results of two linear regressions. The first regresses the BSRT score on the cluster label, age, and sex, while the second regresses BSRT score on AHI, age, and sex. Sex and age are found to be significant in both models, with higher age and being male associated with poorer performance. In model 1, cluster 1 has a significantly better score than reference cluster 4.
@@
However, despite the variation in AHI between the groups and the fact that group 4 has by far the highest AHI, in model 2 the coefficient for AHI is not significant at a significance level of 0.05, suggesting that there may be more to the story when relating OSA severity to a cognitive outcome like BSRT. One aspect of this is that while cluster 1 does not have the lowest AHI, it does have the lowest average $\phi_{R}$, suggesting that the rate of events may matter more in REM sleep than non-REM. Note that group 3 has the lowest AHI and the lowest average $\phi_{N}$, but does not have significantly better performance than group 4 in model 1. Similarly, while cluster 2 also has substantially lower AHI than cluster 4, it also does not differ significantly in BSRT, and its coefficient has a negative point estimate. One interpretation of these results is that the higher AHI of cluster 4 is balanced by cluster 2’s more highly disrupted sleep, which suggests that examining OSA event occurrences and sleep dynamics together as our model does is important for understanding drivers of symptom severity.
@@ 5.5 Alternate clustering based on simple summary statistics
Finally, we examine the differences in the BSRT scores among these alternate groups. Setting cluster 4 as the reference, we find that after controlling for sex and age, patients in cluster 1 have a significantly higher expected BSRT score at a significance level of $0.05$. A full summary of the linear regression coefficients is included in the Supplementary Information. Taken together, the above results suggest that in practice, simple summary statistics such as time in each sleep stage and stage-specific AHI can help give complementary insights to those generated by our model. More broadly, examining how other variables differ by the clusters generated by our model’s random effects can help reveal other variables that could be key to better understanding sleep differences among OSA patients.
@@ 6 Discussion
In this work, we investigated the heterogeneity in sleep stage dynamics and stage-specific rates of apnea and hypopnea events in OSA patients by analyzing data from the APPLES study, which generated several novel insights. We found that for the average patient, having an OSA event during a given epoch increases the probability of awakening or changing sleep stages in the following epoch, although both the baseline probabilities and the magnitude of this change vary substantially among patients. Also, for the average patient, the expected wait time between OSA events is substantially shorter during REM sleep than during non-REM, although again there is significant patient variation in rates. By modeling the covariance of our model’s random effects, we revealed that during REM sleep, patient baseline rates of transitions to Awake or non-REM are positively correlated with one another, as well as being correlated with greater impact of OSA events on these transitions, while during non-REM sleep, baseline rates of transitions to REM or Awake are negatively correlated with the corresponding shifts due to events. Finally, we clustered patients into four interpretable groups based on their random effects. This revealed that other demographic and health variables, such as age, sex, and BMI, vary significantly by event rates and sleep dynamics. Turning to the cognitive performance measure BSRT, we determined that relative to cluster 4, which had the highest average AHI, cluster 1 had significantly better BSRT scores after correcting for age and sex. This result was noteworthy because, while cluster 1 had the lowest rate of events during REM sleep, it did not have the lowest overall AHI, supporting the hypothesis that events in REM sleep may be particularly harmful to cognitive outcomes. In contrast to our clustering approach, a regression of BSRT on AHI directly did not yield a significant association after correcting for sex and age, further emphasizing the limitations of an AHI-based analysis.